Bibliography

home 23/03-04 16:20

[chalmers1982a] A. F. Chalmers. What is this thing called Science? Open University Press, 2

edition, 1982. [hacking1983a] Ian Hacking. Representing and Interventing — Introductory Topics in the Philos-

ophy of Natural Science. Cambridge University Press, 1983. [negroponte1995a] Nicholas Negroponte. Leve digitalt. Tiden Norsk Forlag, 1995. [stoll1995a] Clifford Stoll. Silicon Snake Oil. Macmillian, 1995. [wright1991a] Georg Henrik von Wright. Vitenskapen og fornuften. Cappelens upopulære skrifter.

Cappelen, 1991.